OpenStudy (anonymous):

Solve: -4 (x+1) - 3 = -3 (x-4)

OpenStudy (austinl):

Distribute the -4 and the -3 and type in what you get.

OpenStudy (anonymous):

ok hold on justtt a sec..

OpenStudy (anonymous):

what do you mean distribute?

OpenStudy (austinl):

\[-4 (x+1) - 3 = -3 (x-4)\] You multiply the -4 and -3 through the parenthesis. \[-4x-4-3=-3x+12\] Simplify and get the values with variables on one side of the equation and the numerical values on the other.

OpenStudy (anonymous):

-19 thank you